# DBellQuant: Breaking the Bell with Double-Bell Transformation for LLMs Post Training Binarization

### v1

@@ 4.1 Settings
All experimental procedures were executed utilizing the PyTorch paszke2019huggingface framework in conjunction with the Huggingface library paszke2019huggingface . Models with parameters smaller than 8B are running on a single NVIDIA A30 GPU equipped with 24GB of memory, others are running on a single NVIDIA A100 GPU equipped with 80GB of memory. Consistent with methodologies outlined by Frantar et al. frantar2023gptq and Huang et al. huang2024billm , a calibration dataset comprising 128 samples sourced from the C4 collection raffel2020exploring was employed.
@@
Comprehensive evaluations were carried out across several large language model families, including LLaMA, LLaMA-2, and LLaMA-3 touvron2023llama and the OPT series zhang2022opt . The efficacy of the developed DBellQuant was assessed by calculating the perplexity of the models’ generated text on standard benchmarks: WikiText2 merity2017pointer , and a subset of the C4 data raffel2020exploring . Furthermore, the models’ performance was evaluated based on accuracy across seven zero-shot question-answering tasks: ARC-c clark2018think , ARC-e clark2018think , Hellaswag zellers2019hellaswag , PIQA bisk2020piqa , and Winogrande sakaguchi2020winogrande .
@@
The primary benchmark for comparison for our DBellQuant approach is BiLLM huang2024billm , which represents the current baseline PTQ technique for binary large language models. Additionally, we include other contemporary PTQ algorithms in our comparison, namely Round-to-Nearest (RTN), GPTQ frantar2023gptq , PB-LLM shang2023pb and the current state-of-the-art PTQ technique ARB-LLM li2024arbllmalternatingrefinedbinarizations .
@@ 4.2 Main Results
Table 2: Accuracy of PIQA, ARC-e, ARC-c, HellaSwag, Winogrande and average accuracy of all datasets with BiLLM and our methods on OPT and LLaMA family.
@@
DBellQuant 6 58.00 37.63 18.77 31.83 51.92 39.64
4.3 Ablation Experiments
Optimization Objective
In the definition of DTNP, we employed the L1 loss as described in Eq. 6. To evaluate the effectiveness of the L1 loss, we tested by replacing it with L2 loss and do the same experiments. We observed that the overall performance was inferior to that achieved by L1 loss. Therefore, we selected L1 loss as the loss function for DTNP. Additionally, during the activation-aware initialization process, the parameter $\epsilon$ is introduced as a hyperparameter. We conducted experiments with different values of $\epsilon$, and the results demonstrated that all tested values of $\epsilon$ consistently outperformed prior algorithms. For LLaMA-2-7B, setting $\epsilon$ to 0.85 yielded the best performance. Results are shown in Tab. 4.
Impact of Block Size
We investigated the impact of block size on the quantization performance of DBellQuant by experimenting with block sizes ranging from 64 to 256 columns. The results are shown in Tab. 3. Consistent with other PTQ methods, smaller block sizes yield lower perplexity, but it will increase the diversity of quantization results also increase the weighting overhead. In all experiments shown before we set the block size to 128 because we think it can better balance the bit-width and quantization effect. Furthermore, our method consistently outperforms the current baseline approaches across various block sizes, demonstrating its robustness and generalizability under different configurations.
@@ 4.3 Ablation Experiments
Activation Bit-width Comparisons
We compare the results across different models with different activation bit-widths shown in Fig. 7. Specifically, we observe that when activations are quantized to 8 bits, the model’s perplexity remains nearly unchanged and, in some cases, even decreases. This demonstrates that our approach alleviates the challenges associated with activation quantization. Furthermore, we found that even under lower-bit quantization, such as 6-bit, our model is able to largely maintain its original performance, with only a minimal increase in perplexity. It is noteworthy that for large-scale models such as LLaMA-2 13B and 70B, the perplexity degradation is almost negligible, highlighting the effectiveness and efficiency of our approach in models with substantial parameters.
Method Avtivation Bits Block Size WikiText2 $\downarrow$ C4 $\downarrow$
BiLLM 16 64 20.12 24.46
DBellQuant 8 64 13.67 16.99
DBellQuant 6 64 15.65 19.71
[…]
DBellQuant 8 128 18.65 23.02
DBellQuant 6 128 21.69 25.91
[…]
DBellQuant 8 256 22.34 23.37
DBellQuant 6 256 24.68 28.52
Table 3: Performance of different block size.
@@
Method Loss Function Avtivation Bits WikiText2 $\downarrow$ C4 $\downarrow$
@@
(a) Performance of different loss function.
@@
(b) Performance of different $\epsilon$
Table 4: Ablation study on LLaMA-2-7B,results are measured by perplexity.
5 Conclusion
In this work, we propose DBellQuant, an efficient PTQ method that enables simultaneous weight binarization and activation quantization for LLMs. By analyzing weight distributions conducive to binarization, we design a novel and efficient Learnable Transformation for Dual-Bell algorithm. Specifically, we introduce two customized loss functions and an early stopping mechanism to achieve the desired dual-bell transformation. Furthermore, we analyze and demonstrate how this transformation improves activation quantization, allowing us to achieve near 1-bit weight compression and 6-bit activation quantization with minimal performance degradation—marking the first time such results have been achieved under a PTQ framework. Experiments on open-source LLM families show that DBellQuant significantly advances the performance frontier of SOTA binary PTQ methods.
Appendix A Appendix
A.1 Limitations
Currently, our work only supports quantizing activations to 6 bits. When attempting to quantize activations to lower bit-widths, the model collapses, resulting in a significant drop in performance. However, some recent studies have demonstrated the ability to quantize activations to 4 bits while maintaining competitive model performance. Inspired by these advancements, we aim to adopt similar approaches to further optimize our models. Specifically, we seek to not only binarize weights but also quantize activations to even lower bit-widths, enabling easier deployment and faster computation.
A.2 Broader Impacts
Our work demonstrates the feasibility of simultaneously binarizing weights and quantizing activations to 6 bits in LLMs while maintaining competitive performance. This approach significantly reduces the computational and memory overhead associated with LLM deployment, making them more accessible for resource-constrained environments. By enabling efficient inference, our method contributes to the democratization of advanced AI technologies, reducing the environmental impact of large-scale model deployment. Furthermore, it opens new avenues for research in ultra-low-bit quantization, fostering innovation in model efficiency and scalability.
A.3 Proof for Theorem. 1

### v2

@@ 4.1 Settings
All experimental procedures were executed utilizing the PyTorch (paszke2019huggingface) framework in conjunction with the Huggingface library (paszke2019huggingface). Models with parameters smaller than 8B are running on a single NVIDIA A30 GPU equipped with 24GB of memory, others are running on a single NVIDIA A100 GPU equipped with 80GB of memory. Consistent with methodologies outlined by Frantar et al. (frantar2023gptq) and Huang et al. (huang2024billm), a calibration dataset comprising 128 samples sourced from the C4 collection (raffel2020exploring) was employed.
@@
Comprehensive evaluations were carried out across several large language model families, including LLaMA, LLaMA-2, and LLaMA-3 (touvron2023llama), OPT series (zhang2022opt), Qwen series (qwen2025qwen25technicalreport) and Deepseek Model (Guo_2025). The efficacy of the developed DBellQuant was assessed by calculating the perplexity of the models’ generated text on standard benchmarks: WikiText2 (merity2017pointer), and a subset of the C4 data (raffel2020exploring). Furthermore, the models’ performance was evaluated based on accuracy across five zero-shot question-answering tasks: ARC-c (clark2018think), ARC-e (clark2018think), Hellaswag (zellers2019hellaswag), PIQA (bisk2020piqa), and Winogrande (sakaguchi2020winogrande).
@@
The primary benchmark for comparison for our DBellQuant approach is BiLLM (huang2024billm), which represents the current baseline PTQ technique for binary large language models. Additionally, we include other contemporary PTQ algorithms in our comparison, namely Round-to-Nearest (RTN), GPTQ (frantar2023gptq), PB-LLM (shang2023pb) and the current state-of-the-art PTQ technique ARB-LLM (li2024arbllmalternatingrefinedbinarizations).
Bit-width LTDB Algorithm WikiText2 $\downarrow$ C4 $\downarrow$
16 24.41 27.69
16 17.91 21.83
8 27.39 30.91
8 18.65 23.80
6 32.74 42.48
6 21.69 30.24
4 34.62 45.93
4 23.04 29.90
Table 1: Performance w/o LTDB Algorithm.
Method MathQA LogiQA2
fp16 44.96 $\pm$ 0.91 29.58 $\pm$ 1.15
GPTQ 2bit 20.51 $\pm$ 0.82 14.65 $\pm$ 1.53
BiLLM 1.1bit 24.99 $\pm$ 0.79 20.42 $\pm$ 1.87
DBellQuant 1.1bit 27.65 $\pm$ 0.75 22.78 $\pm$ 1.48
Table 2: DeepSeek-R1-Distill-Qwen-7B results.
Method TextVQA ChartQA MME-Per
fp16 84.9 87.3 1695
GPTQ 2bit 0 0 319
BiLLM 1.1bit 26.3 3.7 638
DBellQuant 1.1bit 28.4 5.2 742
Table 3: Qwen-2.5-VL-7B results.
@@ 4.2 Main Results
Method wikitext2 C4
BiLLM 38.42 40.81
DBellQuant 27.05 31.02
Table 5: Qwen-2-7B results.
Method wikitext2 C4
BiLLM 41.74 53.07
DBellQuant 26.78 30.76
Table 6: Qwen-2.5-7B results.
Model 4096×4096 4096×11008 11008×4096
FP16 0.79463 1.73942 1.82653
BiLLM 0.36842 0.38744 0.43906
ARB-LLMX 0.33180 0.35539 0.36792
DBellQuant 0.27694 0.30085 0.31860
Table 7: Actual inference speed comparison on various linear layers in LLaMA-2-7B.
Table 8: Accuracy of PIQA, ARC-e, ARC-c, HellaSwag, Winogrande and average accuracy of all datasets with BiLLM and our methods on OPT and LLaMA family. Values marked with * indicate usage of NVFP4 format, while other values are represented in integer format.
@@
BiLLM 6 56.09 33.92 16.72 31.75 51.85 38.06
DBellQuant 6 58.00 37.63 18.77 31.83 51.92 39.64
DBellQuant 4* 60.07 40.40 19.89 31.92 53.43 41.14
4.3 Ablation Experiments
Effectiveness of LTDB algorithm
To validate the effectiveness of our advanced LTDB algorithm, we compare them with the vanilla initialization without LTDB. We can observe that under different bitwidths of activation quantization, the models utilizing the proposed LTDB method consistently outperform those relying on vanilla initialization as shown in Tab. 1. This demonstrates that the proposed LTDB method can effectively reduce quantization errors.
Optimization Objective
In DTNP, we adopted L1 loss as the objective function, as it consistently outperformed L2 loss in our experiments. Additionally, we introduced the hyperparameter $\epsilon$ during the activation-aware initialization process and found that all tested values surpassed prior algorithms, with $\epsilon=0.85$ achieving the best performance for LLaMA-2-7B as shown in Tab. 21 and Tab. 16. Detailed parameter settings are provided in the Appendix. A.6.
Impact of Block Size
We evaluated the impact of block size on DBellQuant’s quantization performance using block sizes ranging from 64 to 256 columns, as shown in Tab. 19. Similar to other PTQ methods, smaller block sizes achieve lower perplexity but increase quantization diversity and weighting overhead. In previous experiments, we used a block size of 128, as it balances bit-width and quantization performance effectively. Notably, our method consistently outperforms baseline approaches across all block sizes, highlighting its robustness and generalizability.
Activation Bit-width Comparisons
We compared model performance across different activation bit-widths, as shown in Fig. 7. When activations are quantized to 8 bits, perplexity remains nearly unchanged or even improves. Even with lower-bit quantization, such as 6-bit, the performance remains largely intact, with minimal perplexity increases. Notably, we also conducted experiments with NVFP4 format 4-bit activation quantization (chen2025int) in Tab. 8 and Tab. 9, we can see the model can still retain a significant level of performance. In most cases, it even outperforms 16-bit BiLLM. These results highlight the potential of our method in enabling efficient and reliable model performance under lower-bit settings.
Implementation on Various Advanced Models
To demonstrate the adaptability of DBellQuant beyond standard benchmarks, we extended our evaluation to the Qwen family, reasoning, and multi-modal models. As shown in Tab.5 and Tab.6, DBellQuant consistently outperforms BiLLM on both Qwen-2-7B and Qwen-2.5-7B. Furthermore, we evaluated the reasoning model DeepSeek-R1-Distill-Qwen-7B (Tab.2) and the multi-modal Qwen-2.5-VL-7B (Tab.3). Despite the significant challenge of near 1-bit quantization causing noticeable performance drops, our method still outperforms BiLLM across MathQA, LogiQA2, and visual benchmarks, highlighting its robustness in ultra-low-bit scenarios.
4.4 Time and Memory Analysis
Time Comparison
@@ 4.4 Time and Memory Analysis
Memory Comparison
We include a detailed comparison of bitwidth and model size across different settings. Our method reduces the model size to approximately 1/7 to 1/6 of the original FP16 model, with only a minimal increase in perplexity. Compared to other ultra-low-bit PTQ methods such as BiLLM and PB-LLM, our approach achieves comparable bitwidth and model size, while yielding significantly lower perplexity as shown in Tab. 23. More results can be seen in Appendix. A.12.
5 Conclusion
We propose DBellQuant, an efficient PTQ method enabling simultaneous weight binarization and activation quantization for LLMs. By leveraging weight distributions suited for binarization, we design the Learnable Transformation for Dual-Bell algorithm. This includes two customized loss functions and an early stopping mechanism to achieve the dual-bell transformation. This transformation enhances activation quantization at the same time, enabling near 1-bit weight compression and 6-bit or even 4-bit activation quantization with minimal performance loss—achieving this milestone for the first time in the PTQ framework. Experiments on open-source LLMs across different families and sizes demonstrate that DBellQuant significantly advances the performance of previous SOTA binary PTQ methods.
Impact Statement
This paper enhances inference-time and memory efficiency for LLMs by introducing DBellQuant, a novel post-training quantization framework. DBellQuant significantly reduces computational and memory demands through nearly 1-bit weight compression and 6-bit or even 4-bit activation quantization while maintaining model performance. Its contributions improve the practicality and accessibility of deploying LLMs in resource-constrained environments, such as edge devices. The societal and ethical implications align with standard advancements in ML efficiency, promoting sustainability and usability without introducing new model capabilities, data sources, or application domains. As such, it raises no additional ethical concerns beyond those already associated with LLMs and presents no new broader impacts.
Appendix A Appendix
A.1 Use of Large Language Models
In preparing this manuscript, we exclusively used large language models—GPT-5 (openai2025gpt5)—to refine grammar, flow, and tone at the sentence and paragraph levels. These tools were not used to generate ideas, design experiments, or draw conclusions. All technical content, methodologies, and interpretations were independently authored, rigorously verified, and approved by the authors. To prevent factual inaccuracies or citation errors, every model-edited sentence was reviewed by the authors, and all references were meticulously cross-checked with their original sources. The authors take full responsibility for the accuracy and integrity of this work.
A.2 Limitations
Our current work supports quantizing activations to 6 bits and can handle certain tasks with 4-bit activation quantization while maintaining competitive performance. However, for more challenging tasks, the model may still encounter performance degradation, indicating the need for additional techniques to address these issues. Recent studies have not only demonstrated the feasibility of 4-bit activation quantization but also explored methods to mitigate its limitations. Inspired by these advancements, we aim to further optimize our models by combining low-bit quantization with other techniques.
In addition, sparse attention has recently gained significant attention, with numerous studies highlighting its potential to accelerate inference (wang2025viso; li2025efficient). A key objective of our work is to explore the integration of sparse attention with lower-bit quantization techniques, enabling faster computation while maintaining model performance. This combination could open new possibilities for efficient and practical deployment in real-world scenarios.
A.3 Broader Impacts
Our work demonstrates the feasibility of simultaneously binarizing weights and quantizing activations to 6 bits in LLMs while maintaining competitive performance. This approach significantly reduces the computational and memory overhead associated with LLM deployment, making them more accessible for resource-constrained environments. By enabling efficient inference (yang2024conditional), our method contributes to the democratization of advanced AI technologies, reducing the environmental impact of large-scale model deployment. Furthermore, it opens new avenues for research in ultra-low-bit quantization, fostering innovation in model efficiency and scalability.
A.4 Conceptual Comparison with Prior PTQ Methods
In this section, we clarify the conceptual differences between DBellQuant and several closely related post-training quantization methods, such as SmoothQuant (xiao2023smoothquant), OmniQuant (shao2023omniquant), and QuaRot (ashkboos2024quarot). While all these methods can be interpreted as applying some form of pre/post-transform to weights and/or activations, their design goals, operating bit-widths, and transformation mechanisms are fundamentally different.
Compared to these prior approaches, DBellQuant is not merely another form of scale redistribution:
  • Distribution-shaping for binary PTQ. SmoothQuant and OmniQuant primarily act on the dynamic range of activations and weights, aiming to make low-bit quantization (typically 4–8 bits) more robust by redistributing scaling factors or applying nonlinear rescaling. In contrast, DBellQuant is designed specifically for near 1-bit weight quantization. Our learnable per-channel transform $T$ is optimized with a dedicated objective (DTMD + DTNP) that explicitly drives each channel’s weight distribution from a unimodal “single-bell” shape toward a bimodal “double-bell” shape that is theoretically more compatible with sign-based binarization.
  • Coupled design of weight binarization and activation smoothing. QuaRot applies orthogonal rotations (Hadamard + permutation) to reduce correlations and outliers in weights and activations, which is highly effective for low-bit fixed-point GEMM but is not specifically tailored to the peculiarities of sign-binarized weights. DBellQuant, on the other hand, jointly designs (i) a learnable per-channel scaling of weights $W\odot T$ and (ii) its inverse on activations $X\odot T^{-1}$ so that:
    1. 1.
      the transformed weights become bimodal and thus incur smaller sign-binarization error; and
    2. 2.
      channels with large activation outliers are automatically assigned larger $T_{j}$, so that $T_{j}^{-1}$ suppresses those outliers when applied to activations, making 6-bit activation quantization feasible under a binary weight regime.
  • Binary-centric loss design. While existing methods typically optimize reconstruction error or rely on standard calibration losses, DBellQuant’s LTDB algorithm is explicitly constructed around binary quantization. The Dual-Target Minimum Deviation (DTMD) and Dual-Target Normalized Proportional (DTNP) losses are defined with respect to the two binarization centers derived from the sign function, and the optimization of $T$ is directly driven by the goal of minimizing sign-binarization error. In other words, both the transformation and its learning objective are co-designed for binary PTQ, rather than for generic low-bit quantization.
To summarize, although DBellQuant also employs per-channel transformations like SmoothQuant and interacts with activation outliers like QuaRot, its core contribution is conceptually different: we use a lightweight, equivalence-preserving, learnable transform to reshape the full-precision weight distributions into a bimodal form that is inherently more compatible with sign-based binarization, while the inverse transform automatically smooths activation outliers. This binary-centric distribution-shaping is what enables DBellQuant to achieve near 1-bit weights together with 6-bit activations under a purely post-training setting.
A.5 Effectiveness of DTMD and DTNP Loss
With DTMD loss, T tends to shrink toward zero, leading to a "loss hack" effect that prevents the desired transformation. To address this, we introduce the DTNP loss to ensure a smooth transition. However, using DTNP alone creates a new issue, as its convergence direction can conflict with DTMD. As shown in Tab. 21, using DTNP alone drives the model toward a direction misaligned with DTMD, resulting in degraded performance.
A.6 Hyperparameter setting in training
Our method involves a few key hyperparameters, including the loss coefficient , the number of training epochs, and the learning rate. These were initially determined by validating on the LLaMA2-7B model, where we used 100 for both loss coefficients, 200 epochs, and a learning rate of 0.01.
For the smooth parameter $\epsilon$ , we followed the strategy of SmoothQuant, testing values of 0.75, 0.80, 0.85, and 0.90. The differences in performance were marginal, with 0.85 yielding the best results, and thus selected as the default. The stopping criterion is based on LTDB loss mentioned in Section. 3.3.
Importantly, we found that these hyperparameters generalize well across different architectures and scales—including both the LLaMA and OPT model families—without requiring re-tuning. This demonstrates that our method is robust and not sensitive to hyperparameter choices.
A.7 Visualization of the transformation of weight distributions across different layers
The visualization of weights distribution before and after our proposed method DBellQuant is shown in Fig. 6.
@@ A.7 Visualization of the transformation of weight distributions across different layers
A.8 Detailed application of Transformation
To preserve computational consistency and avoid introducing additional parameters, we apply the transformation $T$ between the LayerNorm and the q/k/v projection layers, as well as between the LayerNorm and the up/gate projection layers. This placement mirrors that of SmoothQuant, ensuring the transformed inputs are compatible with quantization while maintaining the model’s architecture.
For o_projo\_proj, although it is theoretically possible to insert a transformation $T$ between v_projv\_proj and o_projo\_proj to complete the symmetry, we found that this leads to conflicting transformations on v_projv\_proj, degrading performance. Therefore, we do not apply any transformation to o_projo\_proj in practice.
Similarly, for down_projdown\_proj, the presence of activation functions (e.g., GeLU) makes it non-trivial to apply transformations without disrupting the computation graph or introducing inconsistencies. As a result, we do not insert transformations in this part either.
But we need to clarify that both o_projo\_proj and down_projdown\_proj’s weight are binarized in DBellQuant. To clarify, for these two layers, o_projo\_proj and down_projdown\_proj, we did not learn a T-transformation to modify their weights. However, we still applied the original binarization in BiLLM to process these two layers, making our approach a w1a6 method.
A.9 Performance of different activation bit-widths
The performance of different activation bit-widths across different models is shown in Fig. 7.
@@ A.9 Performance of different activation bit-widths
A.10 Results about 4-bit activation on Wikitext2 and C4.
With the help of NVFP4 format on activation quantization, we have conducted experiments on Llama-2-7B to test the perplexity on Wikitext2 and C4 dataset, with nearly 1-bit weight and 4-bit activation. Results can be seen in Tab 9.
Method Bit-width Wikitext2 C4
fp16 16 5.47 7.26
BiLLM 16 32.48 39.38
DBellQuant 8 18.65 23.80
DBellQuant 6 21.69 30.24
DBellQuant 4 23.04 29.90
Table 9: 4-bit activation quantization results.
A.11 Results about training of different methods
Detailed training time and perplexity of different methods across various models are shown in Tab. 10.
Table 10: Training Time and Performance Comparison
OPT-1.3B OPT-2.3B OPT-6.7B LLaMA-1-7B LLaMA-2-7B
mins ppl mins ppl mins ppl mins ppl mins ppl
BiLLM 7 69.97 13 49.55 34 35.36 35 35.04 37 32.48
ARB-LLMX 88 21.81
DBellQuant 9 43.42 17 31.47 45 18.89 47 15.34 49 17.91
A.12 Results about bit-width, model size and perplexity of different methods
Detailed bit-width, model size and perplexity of different methods across various models are shown in Tab. 11, Tab. 12 and Tab. 13.
Table 11: Bit-width, Model Size and Performance Comparison for OPT-6.7B
Methods bit-width model size ppl
FP16 16 12.5GB 10.86
RTN 2 2.15GB 2e4
GPTQ 2 2.15GB 50.19
PB-LLM 1.7 1.95GB 105.16
BiLLM 1.11 1.84GB 35.36
DBellQuant 1.18 1.89GB 18.89
Table 12: Bit-width, Model Size and Performance Comparison for LLaMA-1-7B
Methods bit-width model size ppl
FP16 16 13.5GB 5.68
RTN 2 2.25GB 1e5
GPTQ 2 2.25GB 152.31
PB-LLM 1.7 2.05GB 102.36
BiLLM 1.09 1.97GB 35.04
DBellQuant 1.15 2.02GB 15.34
Table 13: Bit-width, Model Size and Performance Comparison for LLaMA-2-7B
Methods bit-width model size ppl
FP16 16 13.5GB 5.47
RTN 2 2.31GB 1e5
GPTQ 2 2.31GB 60.45
PB-LLM 1.7 2.08GB 69.20
BiLLM 1.08 1.98GB 32.48
DBellQuant 1.15 2.04GB 17.91
A.13 Results about Computation Speedup
To quantify computational speedup, we follow the benchmarking strategy of ARB-LLM and utilize the BitBLAS codebase, which supports mixed-precision GEMM operations for low-bit weights. We benchmark the latency (in milliseconds) of linear layers in LLaMA2-7B using input sequences of length 2048. The results are shown in Tab. 14. Here, we use INT8 format for activation quantization. Our findings are as follows:
1)Weight binarization significantly reduces inference latency compared to FP16 models, thanks to faster memory access and bitwise computation.
2)DBellQuant supports 8-bit and even 4-bit activation using NVFP4 format quantization, enabling efficient inference. This not only accelerates runtime but also yields better perplexity compared to prior binary weight methods like BiLLM and ARB-LLMX.
Table 14: Computation Speed Comparison
Model 4096×4096 4096×11008 11008×4096
FP16 0.79463 1.73942 1.82653
BiLLM 0.36842 0.38744 0.43906
ARB-LLMX 0.33180 0.35539 0.36792
DBellQuant 0.27694 0.30085 0.31860
A.14 Comparison and Combination with Rotation-based Methods
Rotation-based PTQ methods, such as QuaRot with Hadamard-based transforms, effectively reduce outliers in weights and activations. We compared low-bit quantization using these methods alone and in combination with our approach. As shown in Tab.17 and Tab.18, rotation-based PTQ methods experience significant performance degradation when weight precision is reduced below 4-bit. Therefore, we focus on comparisons with state-of-the-art methods for near-1-bit weight quantization. Additionally, we evaluated the combination of DBellQuant and Hadamard-based transforms, such as QuaRot, on the Llama-2-7B model, but this combination performed worse than DBellQuant alone, as shown in Tab. 20. A possible explanation is that DBellQuant transforms the original unimodal distribution into a binarization-friendly bimodal distribution, while Hadamard-based transforms disrupt this structure, negatively impacting quantization performance.
Method Loss Function Activation Bits WikiText2 $\downarrow$ C4 $\downarrow$
@@ A.14 Comparison and Combination with Rotation-based Methods
Table 15: Performance of different loss function
@@
Table 16: Performance of different $\epsilon$
Method 2 bit 1.1 bit
QuaRot+RTN inf inf
QuaRot+GPTQ 22.07 Inf
DBellQuant - 17.91
Table 17: Llama-2-7B results.
Method 2 bit 1.1 bit
QuaRot+RTN inf inf
QuaRot+GPTQ 10.41 Inf
DBellQuant - 12.79
Table 18: Llama-2-13B results.
Method Activation Bits Block Size WikiText2 $\downarrow$ C4 $\downarrow$
BiLLM 16 64 20.12 24.46
DBellQuant 8 64 13.67 16.99
DBellQuant 6 64 15.65 19.71
BiLLM 16 128 32.48 40.52
DBellQuant 8 128 18.65 23.02
DBellQuant 6 128 21.69 25.91
BiLLM 16 256 43.69 43.21
DBellQuant 8 256 22.34 23.37
DBellQuant 6 256 24.68 28.52
Table 19: Performance of different block size.
Method wikitext2 C4
DBellQuant 17.91 21.83
DBellQuant+QuaRot 62.37 76.59
Table 20: Combination results
Loss Function wikitext2 C4
DTMD+DTNP 17.91 21.83
only DTNP 28.74 31.42
Table 21: Loss Function results
Method Model mins $\downarrow$ WikiText2 $\downarrow$
BiLLM LLaMA-1-7B 35 35.04
[…]
DBellQuant LLaMA-1-7B 47 15.34
BiLLM LLaMA-2-7B 37 32.48
DBellQuant LLaMA-2-7B 49 17.91
Table 22: Training time comparison.
Method Bit-width $\downarrow$ Model Size $\downarrow$ Perplexity $\downarrow$
FP16 16 13.5GB 5.47
RTN 2 2.31GB 1e5
GPTQ 2 2.31GB 60.45
PB-LLM 1.7 2.08GB 69.20
BiLLM 1.08 1.98GB 32.48
DBellQuant 1.15 2.04GB 17.91
Table 23: Bit-width and model size comparison.
A.15 Perplexity on C4 dataset
[…]
Full Precision 16 16.07 14.34 12.71 7.34 07.26 06.73 5.71
RTN 16 9999.56 23492.89 9617.07 194607.78 115058.76 46250.21 314504.09
GPTQ 16 6364.65 6703.36 5576.82 186229.5 67954.04 19303.51 13036.32
PB-LLM 16 168.12 222.15 104.78 76.63 80.69 184.67 NAN
BiLLM 16 64.14 44.77 42.13 46.96 39.38 25.87 17.30
[…]
DBellQuant 16 42.57 32.89 21.78 17.60 21.83 15.14 9.49
BiLLM 8 74.56 61.99 40.91 47.13 40.91 21.45 17.72
DBellQuant 8 44.60 32.52 21.56 18.16 23.80 15.56 9.61
BiLLM 6 7348 13445.21 63.41 61.65 63.41 37.66 19.43
DBellQuant 6 57.14 45.24 23.12 19.80 30.24 17.84 10.12
A.16 Visualization of distribution of activation before and after DBellQuant
A.17 Analysis of relative error of the activation
To further validate our method, we conducted experiments demonstrating its effectiveness in reducing the relative error of activations. Specifically, we randomly sampled 128 data points from the C4 dataset and extracted the $q_{\text{proj}}$ inputs from the LLaMA2-7B model using both BiLLM (without applying the inverse transform on activations) and our proposed DBellQuant (with the inverse transform applied).
We evaluated the relative error using two metrics: the Z-score, as you suggested, and the relative deviation error. The formulas are as follows:
  • Z-score:
    $Z=\frac{x-\mu}{\sigma}$
    where $x$ is the value, $\mu$ is the mean, and $\sigma$ is the standard deviation.
  • Relative Deviation Error:
    $\text{Relative Deviation Error}=\frac{x-u}{u}$
    where $u$ is the mean value of the corresponding row.
These metrics allowed us to comprehensively assess the relative error and validate the robustness of our proposed method. The computation results are shown in Tab. 25.
Z-score Relative Deviation Error
BiLLM 0.1584 33.78
DBellQuant 0.1401 30.72
Table 25: Comparison of Z-score and Relative Deviation Error between BiLLM and DBellQuant.
Our method demonstrates its effectiveness not only in reducing the absolute error of activations but also in minimizing outliers from a relative distance perspective, as verified by DBellQuant.
A.18 Algorithm
The detailed algorithm is shown in Algorithm. 1.
[…]
A.19 Proof for Theorem. 1